EP020A

PI: Porcas
Title: DA240

Session: Oct 1996 (session 3)
Frequency: 6 cm
Mode: B
Obs time (UT): 298 EU 20:30(24/10)-10:30(25/10)
Stations: EbJbNtOnToSmUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 4

Comment


To give feedback on this experiment, click here

Experiment Feedback

Torun (Roman Feiler at Fri Oct 25 07:19:36 EDT 1996) Minor failures
6:23 - 6:57 telescope not tracking, occasionally problems with recorder head positioning

Onsala (A. Polatidis, J. Conway at Fri Oct 25 12:15:53 EDT 1996) Success
No known problems

Effelsberg (D Graham at Tue Oct 29 02:41:40 EST 1996) Failure
Operator entered fs in wrong window, sending receiver into frequency switching mode. oscillator approx 30.625kHz too high 2030-0630. If this figure is exact and constant the data can be recovered

Westerbork (tony foley at Tue Oct 29 09:05:26 EST 1996) Minor failures
Looks Ok, but only decoder channel A could read this at double speed.

Noto (Corrado Trigilio at Wed Oct 30 05:48:16 EST 1996) Minor failures
A communication problem with the bbc03 has been detected at 2:00 UT. It has been resolved at 2:40 UT. No other problems reported. Clear sky, cloudy only during the last hour of observations.

Jodrell (Alastair Gunn at Thu Nov 7 06:52:09 EST 1996) Minor failures
Telescope seemed to stay on 0820+560 instead of moving to DA240 at 02:15 UT scan. Subsequent scans seem okay. Late on source for 06:28 scan. Weather okay. Tape sent to Bonn on 28/10/96 via Woodcroft.

Urumqi (Liu Xiang at Mon Nov 18 08:52:40 EST 1996) Minor failures
May be recorded overlaped or not be recorded, please to see the logfile where we made a mistake in the PRCfile (see define TAPEFORM). Clock offset 0.1+/-0.4us.

Simeiz (Alexandre Volvatch at Tue Nov 19 03:52:18 EST 1996) Failure
Not observed - we had problems with MK-3

MPIfR (Walter Alef at Tue Sep 23 17:07:22 METDST 1997) Minor failures
Crimea did not observe. Urumqi had blank tapes. Effelsberg had LO offset of +30630 HZ, for first 10 hours of the observation, but could be fixed during correlation. Rest O.K.